You play a game with two six-sided dice. If you roll a sum of 3 or 8, you win ₱600. If you roll a sum of 10, you win 

₱400. However, you lose ₱500 for anything else. If you continue to play the game, how much do you expect to 

win or lose in the game?


1
Expert's answer
2022-03-03T04:40:16-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 1+1 & 1+2 & 1+3 & 1+4 & 1+5 & 1+6 \\
 \hdashline
 2 & 2+1 & 2+2 & 2+3 & 2+4 & 2+5 & 2+6 \\
 \hdashline
 3 & 3+1 & 3+2 & 3+3 & 3+4 & 3+5 & 3+6\\
 \hdashline
 4 & 4+1 & 4+2 & 4+3 & 4+4 & 4+5 & 4+6 \\
 \hdashline
 5 & 5+1 & 5+2 & 5+3 & 5+4 & 5+5 & 5+6\\
 \hdashline
 6 & 6+1 & 6+2 & 6+3 & 6+4 & 6+5 & 6+6 \\
 \hdashline
\end{array}"

There are "6^2=36" pairs

"1+2=2+1=3"

"2+6=3+5=4+4=5+3=6+2=8"

"4+6=5+5=6+4=10"

"E(X)=600(7\/36)+400(3\/36)-500(26\/36)"

"=-7600/36\approx-211"

I expect to lose ₱211 in the game.